# Statistical evaluation and optimization of entanglement purification protocols

### v1

@@ Abstract
Quantitative characterization of two-qubit entanglement purification protocols is introduced. Our approach is based on the concurrence and the hit-and-run algorithm applied to the convex set of all two-qubit states. We demonstrate that pioneering protocols are unable to improve the estimated initial average concurrence of almost uniformly sampled density matrices, however, as it is known, they still generate pairs of qubits in a state that is close to a Bell state. We also develop a more efficient protocol and investigate it numerically together with a recent proposal based on an entangling rank-two projector. Furthermore, we present a class of variational purification protocols with continuous parameters and optimize their output concurrence. These optimized algorithms turn out to surpass former proposals and our new protocol by means of not wasting too many entangled states.
@@ I Introduction
Entanglement purification protocols aim to overcome the destructive effects of non-ideal channels by generating highly entangled states from a large number of noisy entangled states [1]. In this approach, the obtained almost perfectly entangled pairs are used in quantum teleportation [2], and thus quantum data can be transmitted across the channels. The other possible solution to this problem is to use quantum error correction [3], when quantum data is sent through the channel by adding enough redundancy, e.g., increasing the number of qubits, such that the original information despite the presence of noise is recoverable. While in quantum error correction, the sources of errors and their models have to be identified, the first entanglement purification protocols [4, 5] offer solutions for general noisy two-qubit states. However, these protocols convert a general two-qubit state to either Werner or Bell diagonal states by using local random transformations, which waste useful entanglement [6, 7]. This was found by investigating particular entangled states, whose entanglement is destroyed in the first step of the protocol, and this issue was also confirmed in a different entanglement purification approach [8]. A quantitative characterization of the ratio between wasted and improved entangled states is missing and this paper is devoted to investigating this question.
@@
The quest for optimality is an ongoing and important research in quantum physics ranging from experimental protocols to algorithms. In most cases, numerical optimization based on continuous and discrete parameters is carried out to enhance performances with respect to different figures of merit. This has recently led to the improvement of entanglement generation [9], unitary compilation and state preparation [10], quantum error correction [11], communication [12, 13], and algorithms [14], which are only a few examples of a considerable literature. Concerning entanglement purification, discrete optimizations have been considered for Werner states [15] and we have started to investigate continuous optimization for certain families of states [16]. Therefore, in this paper, we not only evaluate some existing proposals but also search for more optimal protocols with less wasted entangled states.
@@
In this study, we avail ourselves of the hit-and-run algorithm, which generates asymptotically and effectively uniformly distributed two-qubit density matrices [17]. It is known that approximately $24.24$ of the generated two-qubit states are separable and this numerical result has also been confirmed by other methods [18, 19, 20, 21]. Therefore, almost one-fourth of the states are immediately useless in an entanglement purification protocol. In the case of entangled states, we use the concurrence [22] to measure the improvement or deterioration induced by a protocol. We calculate the average concurrence over the whole sample, which will be our cost function. Any change of this cost function characterizes only the whole sample of two-qubit states, whereas the concurrence of individual states may show different behaviors. The obtained estimates allow us to compare quantitatively some existing entanglement purification protocols and to search numerically for more optimal scenarios. These scenarios consist of the use of the $SU(4)\times SU(4)$ group to find optimal locally entangling gates. We assume throughout the whole paper that these operations can be performed without errors on both sides of the noisy channel.
@@
Our search for optimal protocols is based on the limited memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS-B) optimization algorithm [23, 24, 25] with bounds. To use this method we employ the Euler angle parametrization of the $SU(4)\times SU(4)$ group [26], where the group is constructed from angles, which form a hyperrectangle in a $30$-dimensional Euclidean vector spaces. Thus, every vector represents an element of the group and its neighborhood is defined by the Euclidean norm. The resulting parametrization allows us to construct a variational cost function that depends on a unitary in $SU(4)\times SU(4)$ and that can be minimized with the L-BFGS-B method thanks to automatic differentiation [27]. With this strategy, we demonstrate that it is possible to increase the percentage of entangled states purified by computer-designed protocol and achieve better performances. Furthermore, our approach introduces a systematic and general way to compare different proposals for entanglement purification protocols, where more realistic experimental considerations can be incorporated upon request.
@@
The paper is organized as follows. In Sec. II, we recall the mathematical description of some known entanglement purification protocols. Furthermore, we present a new controlled-NOT(CNOT) gate-based protocol and also our variational approach. A brief description of the numerical approach is given in Sec. III. Numerical results are presented and discussed in Sec. IV. Finally, Sec. V contains our conclusions. Some technical details are provided in the Appendix.
@@ II Entanglement purification
$r_{1}+r_{2}+r_{3}+r_{4}=1,\quad r_{ij}=\left(r_{ji}\right)^{*}.$
In the following, we give a short overview of three purification protocols, introduce a new CNOT-based approach, and our variational method for the search of more optimal strategies.
@@ II.1 Bennett protocol
where $I_{4}$ is the $4\times 4$ identity matrix. The transformation of a general state $\rho$ of Eq. (2) into a Werner state $\rho_{W}$ can be achieved by local random unitary rotations [6], i.e., so-called twirling operation, which is given by
@@
  • Measure the target pair ($A_{2},B_{2}$) in eigenbasis of the Pauli matrix $\sigma_{z}$ with corresponding results $(m,n)$, where $m,n\in\{0,1\}$. Keep the pair ($A_{1},B_{1}$) if the measurement result is either $m=n=0$ or $m=n=1$ and finally perform a $\sigma_{y}$ rotation on $A_{1}$.
@@ II.2 Deutsch protocol
where the $K_{i}$s are given in Eq. (5). The Deutsch protocol can be summarized in three steps:
@@
Applying the purification protocol we obtain a new Bell diagonal state $\rho^{\prime}_{B}$, which is described by the following map:
@@
Depending on which of the above conditions is fulfilled, the protocol distills asymptotically either |$\ket{4}\bra{4}$ | or |$\ket{2}\bra{2}$ | [29], i.e, the Bell states |$\ket{4}$ and |$\ket{2}$.
@@ II.3 Matter-field interaction-based protocol
A key step in the previous protocols is the application of an entangling unitary transformation in the nodes $A$ and $B$. The motivation to choose the abstract CNOT gate comes mainly from classical computer science. However, any entangling quantum operation can serve the same purpose as shown by one of us in a cavity QED setup, where an entangling transformation emerges from matter-field interactions and is modeled by a rank-two projector [30]
@@
$r^{\prime}_{2}$ $=$ $\frac{r^{2}_{2}+r^{2}_{4}-r^{2}_{24}-r^{2}_{42}}{D},\quad r^{\prime}_{4}=2\frac{r_{1}r_{3}-|r_{13}|^{2}}{D},$
@@
where success probability $P_{s}=D/2$ and
@@
This protocol was analysed in Ref. [8] and it was found that either
@@
is fulfilled, then $\rho^{\prime}$ becomes more entangled than $\rho$. In the asymptotic limit, the protocol converts all states fulfilling Eq. (17) into |$\ket{1}\bra{1}$ | and Eq. (18) into |$\ket{2}\bra{2}$ |, respectively.
@@ II.4 A CNOT-based protocol
$r^{\prime}_{3}$ $=$ $2\frac{r_{1}r_{4}-|r_{14}|^{2}}{E},\quad r^{\prime}_{4}=\frac{r^{2}_{1}+r^{2}_{4}+r^{2}_{14}+r^{2}_{41}}{E},$
@@
To increase the degree of entanglement of $\rho^{\prime}$, the state $\rho$ has to fulfill either
@@
is fulfilled. These conditions are more restrictive than their counterparts in Eqs. (21) and (22). They work for Bell diagonal states, but many states do not obey them, and therefore for a general purification strategy the pair ($A_{1},B_{1}$) has to be discarded, whenever the pair ($A_{2},B_{2}$) is measured in the state |$\ket{00}$. If the states fulfill condition given in Eq. (21) then the protocol can distill |$\ket{4}\bra{4}$ |, otherwise in the case of Eq. (22), the state |$\ket{2}\bra{2}$ | is obtained. It is worth noting that these Bell states can also be reached not only in the asymptotic limit. Based on the second example in Ref. [8], one can pick $r_{2}=c$, $r_{1}=r_{4}=(1-c)/2$, and $r_{14}=\left(r_{41}\right)^{*}=i(1-c)/2$ with $c\in(0,1]$ to observe that one iteration of the protocol yields |$\ket{2}\bra{2}$ | with success probability $c^{2}/2$. This state cannot be purified by the Bennett protocol. If $c\in(0.5,1]$, then the Deutsch protocol approaches the Bell state |$\ket{2}$ only asymptotically.
@@ II.5 Variational purification protocols
where the asterisk denotes the complex conjugation in the standard basis. The advantage of the concurrence lies in the fact that it treats all maximally entangled states equivalently. As we have discussed in the previous subsections, the maps there can have two stable fixed points corresponding to two Bell states, and they can approach both of them depending on the properties of the input density matrix. Therefore, picking a fidelity with respect to a Bell state, would not work, because the optimization would suppress the convergence towards another Bell state or any maximally entangled state, which can be reached otherwise. A different argument against the use of fidelity with an example is given in Ref. [16].
The purification protocol for our numerical investigations reads:
@@
    $\varrho^{\prime}=\frac{\Pi\varrho\Pi^{\dagger}}{\mathrm{Tr}\left\{\Pi^{\dagger}\Pi\varrho\right\}}.$ (27)
@@
Let us consider a general unitary matrix $V(\boldsymbol{\alpha}_{AB})\in SU(4)\otimes SU(4)$ acting on $A_{1},A_{2}$ and $B_{1},B_{2}$. We employ the Euler-angle parametrization of $SU(4)$ [26]:
where $\sigma_{i},\ i=1,...,15$ form a Gell-Mann type basis of the Lie group $SU(4)$, see the Appendix in [16]. Then, we have
where $\boldsymbol{\alpha}_{AB}=(\boldsymbol{\alpha}_{A},\boldsymbol{\alpha}_{B})^{T}$ and thus $\boldsymbol{\alpha}_{AB}\in\mathbb{R}^{30}$. If we employ this unitary matrix as the entangling operation, the output density matrix $\rho_{\text{out}}$ of the variational purification protocol will depend on the value of $\boldsymbol{\alpha}_{AB}$:
@@
The Euler-angle parametrization guarantees that the group is not naively overcounted [26]. It is worth noting that due to the nature of the problem, one cannot consider operations that improve entanglement across the channel, i.e., general $SU(16)$-operations on $A$ and $B$, because badly entangled states are the result of the noisy channel. In the numerical simulations, we will specify, when $\Pi$ in Eq. (27) is equal to $V(\boldsymbol{\alpha}_{AB})$ or another matrix.

### v2

@@ Abstract
Quantitative characterization of two-qubit entanglement purification protocols is introduced. Our approach is based on the concurrence and the hit-and-run algorithm applied to the convex set of all two-qubit states. We demonstrate that pioneering protocols are unable to improve the estimated initial average concurrence of almost uniformly sampled density matrices, however, as it is known, they still generate pairs of qubits in a state that is close to a Bell state. We also develop a more efficient protocol and investigate it numerically together with a recent proposal based on an entangling rank-$2$ projector. Furthermore, we present a class of variational purification protocols with continuous parameters and optimize their output concurrence. These optimized algorithms turn out to surpass former proposals and our protocol by means of not wasting too many entangled states.
@@ I Introduction
Entanglement purification protocols aim to overcome the destructive effects of non-ideal channels by generating highly entangled states from a large number of noisy entangled states [1]. In this approach, the almost perfectly entangled pairs obtained are used in quantum teleportation [2], and thus quantum data can be transmitted across the channels. The other possible solution to this problem is to use quantum error correction [3], when quantum data is sent through the channel by adding enough redundancy, e.g., increasing the number of qubits, such that the original information is recoverable even in the presence of noise. While in quantum error correction the sources of errors and their models have to be identified, the first entanglement purification protocols [4, 5] offer solutions for general noisy two-qubit states. However, these protocols convert a general two-qubit state to either Werner or Bell diagonal states by using local random transformations, which waste useful entanglement [6, 7]. This was found by investigating particular entangled states, whose entanglement is destroyed in the first step of the protocol, and this issue was also confirmed in a different entanglement purification approach [8]. A quantitative characterization of the ratio between wasted and improved entangled states is missing and this paper is devoted to investigating this question.
@@
The quest for optimality is ongoing and important research in quantum physics ranging from experimental protocols to algorithms. In most cases, numerical optimization based on continuous and discrete parameters is carried out to enhance performances with respect to different figures of merit. This has recently led to the improvement of entanglement generation [9], unitary compilation and state preparation [10], quantum error correction [11], communication [12, 13], and algorithms [14], which are only a few examples of the vast literature. Concerning entanglement purification, discrete optimizations have been considered for Werner states [15] and we have started to investigate continuous optimization for certain families of states [16]. Therefore, in this paper we not only evaluate some existing proposals but also search for more optimal protocols with a lower amount of wasted entangled states.
@@
In this study, we avail ourselves of the hit-and-run algorithm to generate asymptotically and effectively uniformly distributed two-qubit density matrices [17]. It is known that approximately $24.24$ of the generated two-qubit states are separable, a numerical result that has also been confirmed by other methods [18, 19, 20, 21]. Therefore, almost one-fourth of the states are immediately useless in an entanglement purification protocol. In the case of entangled states, we use the concurrence [22] to measure the improvement or deterioration induced by a protocol. The choice of the concurrence and its advantage over the usually employed fidelity will be explained in the context of those protocols, which have two Bell states as stable fixed points. We calculate the average concurrence over the whole sample, which will be our cost function. Any change of this cost function characterizes only the whole sample of two-qubit states, whereas the concurrence of individual states may show different behaviors. The obtained estimates allow us to compare quantitatively some existing entanglement purification protocols and to search numerically for more optimal scenarios. These scenarios consist of the use of the $SU(4)\times SU(4)$ group to find optimal locally entangling gates. We assume throughout the whole paper that these operations can be performed without errors on both sides of the noisy channel.
@@
Our search for optimal protocols is based on the limited memory Broyden-Fletcher-Goldfarb-Shanno (LBFGSB) optimization algorithm with bounds [23, 24]. To use this method we employ the Euler angle parametrization of the $SU(4)\times SU(4)$ group [25], where the group is constructed from angles, which form a hyperrectangle in a $30$-dimensional Euclidean vector spaces. Thus, every vector represents an element of the group and its neighborhood is defined by the Euclidean norm. The resulting parametrization allows us to construct a variational cost function that depends on a unitary in $SU(4)\times SU(4)$ and that can be minimized with the LBFGSB method due to automatic differentiation [26]. With this strategy, we demonstrate that it is possible to increase the percentage of entangled states purified by computer-designed protocol and achieve better performances. Furthermore, our approach introduces a systematic and general way to compare different proposals for entanglement purification protocols, where more realistic experimental considerations can be incorporated upon request.
@@
The paper is organized as follows. In Sec. II, we recall the mathematical description of some known entanglement purification protocols. Furthermore, we present a controlled-NOT(CNOT) gate-based protocol and also our variational approach. A brief description of the numerical approach is given in Sec. III. Numerical results are presented and discussed in Sec. IV. Section V contains a summary and our conclusions. Some technical details are provided in the Appendices.
@@ II Entanglement purification
$r_{1}+r_{2}+r_{3}+r_{4}=1,\quad r_{ij}=\left(r_{ji}\right)^{*}.$ (3)
In the following, we give a short overview of three purification protocols and introduce a CNOT-based approach and our variational method for the search of more optimal strategies.
@@ II.1 Bennett protocol
where $I_{4}$ is the $4\times 4$ identity matrix. The transformation of a general state $\rho$ of Eq. (2) into a Werner state $\rho_{W}$ can be achieved by local random unitary rotations [6], i.e., the so-called twirling operation, which is given by
@@
  4. (iv)
    Measure the target pair ($A_{2},B_{2}$) in the eigenbasis of the Pauli matrix $\sigma_{z}$ with corresponding results $(m,n)$, where $m,n\in\{0,1\}$. Keep the pair ($A_{1},B_{1}$) if the measurement result is either $m=n=0$ or $m=n=1$ and finally perform a $\sigma_{y}$ rotation on $A_{1}$.
@@ II.2 Deutsch protocol
where the $K_{i}$ are given in Eq. (6). The Deutsch protocol can be summarized in three steps:
  1. (i)
@@
After applying the purification protocol we obtain a new Bell diagonal state $\rho^{\prime}_{B}$, which is described by the map
@@
Depending on which one of the above conditions is fulfilled, the protocol distills asymptotically either |$\ket{4}\bra{4}$ | or |$\ket{2}\bra{2}$ | [28], i.e, the Bell states |$\ket{4}$ and |$\ket{2}$.
@@ II.3 Matter-field interaction-based protocol
A key step in the previous protocols is the application of an entangling unitary transformation in the nodes $A$ and $B$. The motivation to choose the abstract CNOT gate comes mainly from classical computer science. However, any entangling quantum operation can serve the same purpose as shown in Ref. [29] for a cavity QED setup, where an entangling transformation emerges from matter-field interactions and is modeled by a rank-$2$ projector [29]
@@
$r^{\prime}_{2}$ $=$ $\frac{r^{2}_{2}+r^{2}_{4}-r^{2}_{24}-r^{2}_{42}}{D},\quad r^{\prime}_{4}=2\frac{r_{1}r_{3}-|r_{13}|^{2}}{D},$ (16)
@@
$r^{\prime}_{13}$ $=$ $r^{\prime}_{14}=r^{\prime}_{23}=r^{\prime}_{24}=0,$
where the success probability $P_{s}=D/2$ and
@@
Due to the relations in Eq. (3), we also have $r^{\prime}_{21}=\left(r^{\prime}_{12}\right)^{*}$ and $r^{\prime}_{43}=\left(r^{\prime}_{34}\right)^{*}$. Therefore, $r^{\prime}_{1}$, $r^{\prime}_{2}$, $r^{\prime}_{3}$, $r^{\prime}_{4}$, $r^{\prime}_{12}$, $r^{\prime}_{21}$, $r^{\prime}_{34}$, and $r^{\prime}_{43}$ are the nonvanishing elements of $\rho^{\prime}$. This protocol was analysed in Ref. [8] and it was found that either
@@
is fulfilled, then $\rho^{\prime}$ becomes more entangled than $\rho$. In a further iteration, the output state $\rho^{\prime\prime}$ remains in the same form as the input state $\rho^{\prime}$ and according to Eq. (16) its elements read
$r^{\prime\prime}_{1}$ $=$ $\frac{r^{\prime 2}_{1}+r^{\prime 2}_{3}}{D^{\prime}},\quad r^{\prime\prime}_{2}=\frac{r^{\prime 2}_{2}+r^{\prime 2}_{4}}{D^{\prime}},\quad r^{\prime\prime}_{3}=2\frac{r^{\prime}_{2}r^{\prime}_{4}}{D^{\prime}}$
$r^{\prime\prime}_{4}$ $=$ $2\frac{r^{\prime}_{1}r^{\prime}_{3}}{D^{\prime}},\quad r^{\prime\prime}_{12}=\frac{r^{\prime 2}_{12}+r^{\prime 2}_{34}}{D^{\prime}}=\left(r^{\prime\prime}_{21}\right)^{*},$
$r^{\prime\prime}_{34}$ $=$ $2\frac{r^{\prime}_{21}r^{\prime}_{43}}{D^{\prime}}=\left(r^{\prime\prime}_{43}\right)^{*},$
where
$D^{\prime}=(r^{\prime}_{1}+r^{\prime}_{3})^{2}+(r^{\prime}_{2}+r^{\prime}_{4})^{2}.$ (20)
In the asymptotic limit, the protocol converts all states fulfilling Eq. (18) into |$\ket{1}\bra{1}$ | and Eq. (19) into |$\ket{2}\bra{2}$ |.
@@ II.4 A CNOT-based protocol
$r^{\prime}_{3}$ $=$ $2\frac{r_{1}r_{4}-|r_{14}|^{2}}{E},\quad r^{\prime}_{4}=\frac{r^{2}_{1}+r^{2}_{4}+r^{2}_{14}+r^{2}_{41}}{E},$ (21)
@@
According to Eq. (3), we also have $r^{\prime}_{31}=\left(r^{\prime}_{13}\right)^{*}$ and $r^{\prime}_{42}=\left(r^{\prime}_{24}\right)^{*}$. Therefore, $r^{\prime}_{1}$, $r^{\prime}_{2}$, $r^{\prime}_{3}$, $r^{\prime}_{4}$, $r^{\prime}_{13}$, $r^{\prime}_{31}$, $r^{\prime}_{24}$, and $r^{\prime}_{42}$ are the nonvanishing elements of $\rho^{\prime}$. To increase the degree of entanglement of $\rho^{\prime}$, the state $\rho$ has to fulfill either
@@
is fulfilled. These conditions are more restrictive than their counterparts in Eqs. (23) and (24). They work for Bell diagonal states, but many states do not obey them, and therefore for a general purification strategy the pair ($A_{1},B_{1}$) has to be discarded, whenever the pair ($A_{2},B_{2}$) is measured in the state |$\ket{00}$. In a further iteration, the output state $\rho^{\prime\prime}$ remains in the same form as the input state $\rho^{\prime}$ and according to Eq. (21) its elements read
$r^{\prime\prime}_{1}$ $=$ $2\frac{r^{\prime}_{2}r^{\prime}_{3}}{E^{\prime}},\quad r^{\prime\prime}_{2}=\frac{r^{\prime 2}_{2}+r^{\prime 2}_{3}}{E^{\prime}},\quad r^{\prime\prime}_{3}=2\frac{r^{\prime}_{1}r^{\prime}_{4}}{E^{\prime}}$
$r^{\prime\prime}_{4}$ $=$ $\frac{r^{\prime 2}_{1}+r^{\prime 2}_{4}}{E^{\prime}},\quad r^{\prime\prime}_{13}=2\frac{r^{\prime}_{24}r^{\prime}_{31}}{E^{\prime}}=\left(r^{\prime\prime}_{31}\right)^{*},$
$r^{\prime\prime}_{24}$ $=$ $\frac{r^{\prime 2}_{31}+r^{\prime 2}_{24}}{E^{\prime}}=\left(r^{\prime\prime}_{42}\right)^{*},$
where
$E^{\prime}=(r^{\prime}_{1}+r^{\prime}_{4})^{2}+(r^{\prime}_{2}+r^{\prime}_{3})^{2}.$ (27)
If the states fulfill the condition given in Eq. (23) then the protocol can distill |$\ket{4}\bra{4}$ |, otherwise in the case of Eq. (24), the state |$\ket{2}\bra{2}$ | is obtained. It is worth noting that these Bell states can also be reached not only in the asymptotic limit. Based on the second example in Ref. [8], one can pick $r_{2}=c$, $r_{1}=r_{4}=(1-c)/2$, and $r_{14}=\left(r_{41}\right)^{*}=i(1-c)/2$ with $c\in(0,1]$ to observe that one iteration of the protocol yields |$\ket{2}\bra{2}$ | with success probability $c^{2}/2$. This state cannot be purified by the Bennett protocol. If $c\in(0.5,1]$, then the Deutsch protocol approaches the Bell state |$\ket{2}$ only asymptotically.
@@ II.5 Variational purification protocols
where the asterisk denotes the complex conjugation in the standard basis. The advantage of the concurrence lies in the fact that it treats all maximally entangled states equivalently. As we have discussed in the previous sections, the maps there can have two stable fixed points corresponding to two Bell states and they can approach both of them depending on the properties of the input density matrix. Therefore, picking a fidelity with respect to a Bell state would not work, because the optimization would suppress the convergence towards another Bell state or any maximally entangled state, which can be reached otherwise. A different argument against the use of fidelity with an example is given in Ref. [16].
Let us consider a general unitary matrix $V\in SU(4)\otimes SU(4)$ acting on $A_{1},A_{2}$ and $B_{1},B_{2}$. We employ the Euler-angle parametrization of $SU(4)$ [25],
where $\sigma_{i},\ i=1,...,15$ form a Gell-Mann type basis of the Lie group $SU(4)$ (see the Appendix in Ref. [16], where the same approach to angle parametrization was used). Then we have
where $\boldsymbol{\alpha}_{AB}=(\boldsymbol{\alpha}_{A},\boldsymbol{\alpha}_{B})^{T}$ and thus $\boldsymbol{\alpha}_{AB}\in\mathbb{R}^{30}$. The Euler-angle parametrization guarantees that the group is not naively overcounted [25]. It is worth noting that due to the nature of the problem, one cannot consider operations that improve entanglement across the channel, i.e., general $SU(16)$ operations on $A$ and $B$, because badly entangled states are the result of the noisy channel.
The purification protocol for our numerical investigations reads as follows.
  1. (i)
@@
    $\varrho^{\prime}=\frac{\Pi\varrho\Pi^{\dagger}}{\mathrm{Tr}\left\{\Pi^{\dagger}\Pi\varrho\right\}},$ (32)
    where $\Pi$ is a quantum operation. In the numerical simulations, apart from one particular case, $\Pi$ equals $V(\boldsymbol{\alpha}_{AB})$.
  2. (ii)
@@
If we employ $\Pi=V(\boldsymbol{\alpha}_{AB})$ as the entangling operation, the output density matrix $\rho_{\text{out}}$ of the variational purification protocol will depend on the value of $\boldsymbol{\alpha}_{AB}$: